# Tileable Surfaces

### v1

@@ Abstract.
We define a class of $C^{k}$-regular surfaces, $k\geq 1$, tileable surfaces, that admit geometric tilings by a finite number of congruence classes of tiles. We show how to construct many examples, and examine the relationship with the well known tilings of the plane and sphere, as well as monohedral polyhedral surfaces.
@@ 1. Introduction
Tilings of the plane, sphere and hyperbolic space are well known. But if we consider physical manifestations of tiling, that is, a figure or object composed of many copies of the same shape, or a limited number of distinct shapes, there is no special reason why a given tile should be of constant curvature, or that the total space need be homogeneous. In manufacturing and architectural scenarios, there is an economic case for constructing a surface from many copies of a limited number of distinct shapes or tiles. Recently, there has been some serious work related to this problem, in the special case of polyhedral surfaces and variants – see the literature review in Section 2 below.
@@
where $m$ is an extended natural number and,
@@
Even if we allow for a large (finite) number $n$ of prototiles, it is easy to see that there are restrictions on the global shape of a tileable surface. For instance:
@@
To see that this holds, note that, since $S$ is tiled by a finite number of distinct compact prototiles, the curvature must achieve both its maximum and minimum values, and hence be bounded away from zero. By Efimov’s theorem [7], there is no complete $C^{2}$-immersed surface in $\mathbb{R}^{3}$ with curvature bounded above by a negative constant. If the curvature is positive the surface must be compact by the Bonnet-Myers theorem, and hence a topological sphere by the Gauss-Bonnet theorem.
Similarly, given a $C^{2}$ monotiled surface, although it is obviously possible to deform the prototile and apply the corresponding deformation to all the other copies to remain a $C^{2}$ monotiled surface with the same toplogy and face combinatorics, there are nevertheless restrictions on the deformed shape. By the Gauss-Bonnet theorem we have:
@@
    Tileable surfaces with boundary (“incomplete”).
@@ 1.1. Relation between tileable surfaces and tilings of the plane or sphere
Figure 2. From left to right: monohedral polyhedron (biarc-hull [9]); projection to the sphere; more general $C^{1}$ tiling; geometrically distinct tiled surface with the same prototile.
@@ 1.2. Relation between tileable surfaces and monohedral polyhedra
One idea for producing examples of tileable surfaces is to take a monohedral polyhedron and, if possible, deform the model face until the faces fit together tangent continuously along the edges (Figure 2). This is obviously possible in some cases. Conversely, given a tiled surface, it may be possible to choose a small number of well defined “vertices” on the boundary of the prototile, and join them up to form a polygon, such that the polygons fit together to form a polyhedral surface. This can be done for the first surface in Figure 1, to form a tiled plane. For the torus shown in the same figure, the resulting polyhedron collapses because the vertices are all in the same plane, and for some more general tiled surfaces (Figure 4) there are no canonical points to choose as vertices.
If we restrict to finite edge (an analogue of edge to edge tilings) tileable surfaces (see definition below), then we can say that compact finite edge monotileable surfaces correspond to (possibly degenerate) monohedral polyhedra, but not always the other way around (See Theorem 12 below and Figure 3).
A basic problem then is to understand which monohedral polyhedra can be smoothed to become regular tiled surfaces.
@@ 2.1. Doubly and triply periodic surfaces
These are true tileable surfaces, according to our definition. Any 2-dimensional or 3-dimensional lattice structure gives us many examples of monotiled surfaces, simply by placing a surface in the fundamental domain in such a way that translated copies of it are joined tangent continuously. A deformation of a periodic plane-tiling (Figure 1, left) is an example of a doubly periodic surface. More generally, so-called triply periodic surfaces in $\mathbb{R}^{3}$ can easily be constructed by placing a suitable surface inside the fundamental domain of a 3-dimensional lattice. It is possible to arrange that the fundamental piece is also tiled by congruent copies of a topological disc, thus the entire surface satisfies our definition of a monotiled surface. Triply periodic minimal surfaces, first discovered by Schwarz and his student in the 19th century, are well known (see, e.g. [22], [19], [12]) and have recently become very popular in practical applications such as in additive manufacturing and in tissue engineering: see, e.g., [24], [10] and many other recent works.
@@ 2.2. Tilings of constant curvature surfaces
This topic is well known, and extensively studied. See [14], [25], [27] and [1], for an introduction to tilings, and [28] for a survey on monotiling of the plane. Common themes are classifications of patterns, group actions, tilings with as small a number of distinct tiles as possible, non-repeating tilings, tilings with different edge colorings etc. However, we have been unable to find any systematic attempt to extend this study to smooth surfaces of non-constant curvature.
@@ 2.3. Polyhedral surfaces
Steinitz’s theorem states that convex polyhedra are in one-one correspondence with 3-connected planar graphs. However, there are restrictions on the geometry: for instance, not every such graph can be realized as a monohedral polyhedron. Grünbaum [13] shows that there exist triangulated spheres that are not realized by any acoptic (meaning homeomorphic to a sphere) monohedral polyhedron.
A monohedral polyhedron is called an isohedron if all faces lie within the polyhedron’s symmetry orbit, i.e. the rigid motions of the entire polyhedron act transitively on the faces. Examples include the Platonic solids, the Catalan solids, the bipyramids and the trapezohedra. Convex isohedra are classified, with a list of 30 types. Even in the convex case, general monohedral polyhedra do not appear to be classified. Examples that are not isohedra include the triaugmented triangular prism (Figure 3), which has equilateral triangular faces, and the rhombic icosahedron which has 20 rhombic faces. Eppstein [9] gives an infinite family of monohedral polyhedra monotiled by isosceles triangles.
@@ 2.4. Polyhedral and semidiscrete approximations
Eigensatz et al [8] describe the “paneling problem” where a free form surface is to be approximated by as few as possible distinct panels to minimize the number of molds to be used in the construction. Liu et al [20] study a similar problem, using polygonal faces. Singh and Schaeffer [26] give an algorithm where the input is an arbitrary triangulated surface and the output is an approximation using a small number of polygonal faces. Fu et al [11] give an algorithm for optimization of a quadmesh approximation of a surface using a small number of distinct (non-planar) quadrilaterals.
Chen et al [4] consider a similar problem for a shell structure: approximate the surface by a piecewise linear shell structure, where the goal is to have many identical parts. The difference from polyhedral surfaces is the thickness.
Bo et al [3] (see also [2]), replace polyhedral surfaces with circular arc structures, that is surfaces defined by 2D mesh combinatorics, where the edges are circular arcs that meet with a common tangent plane. Their study shares features of what we discuss below, in that the edges of the faces are tangent continuous. However they do not consider surfaces with congruent faces - instead they aim to construct surfaces with congruent vertices.
@@ 2.5. Mosaic methods
Another approximation concept for smooth surfaces, mosaic methods [23], [6] [5],[17], means approximating a 3D surface by covering as much as possible of it with flat tiles drawn from a given collection. Small gaps are allowed between the tiles. The measure of success for a tile set together with an algorithm is to minimize the gap area while closely approximating the surface.
@@ 2.6. Tiling inspired decoration of arbitrary surfaces
Any surface has the conformal structure of one of the standard 3 constant curvature space forms, which allows for texture maps conformally equivalent to the many known tilings of space forms (see, e.g. Gu et al [15]). These are not tilings in the sense discussed here, however, since conformal maps preserve angles but not size.
Jiang et al [18] explore the idea of extending the common tiling patterns on the plane to polyhedral surfaces by another idea. The surface tiles are defined by lifting planar tiles onto paraboloids (approximating the surface around a point by the curvature paraboloid). The method gives a natural way to produce tiling-like patterns on general surfaces, but the tiles are not congruent to each other. The shape of the lift to the surface of a particular model tile will depend on the curvature of the surface at that point.
@@ 3.1. Rigidity of a tile
    There is a unique list $T_{1},T_{2},\dots T_{N}$ that constitutes the set of neighbours of $T$ in any tiled surface containing $T$.
@@
The tiles in Figure 1 are rigid, whilst Figures 2 and 4 show examples of tiles that are not rigid. The tile in Figure 2 can be surrounded by itself in only a finite number of ways. The tile in Figure 4 has a circle as its boundary, and a constant normal along the boundary. There are uncountably many ways to fit two copies together.
@@
    and that each set $\psi_{i}(T_{i_{j}})$ is in fact a tile of $S$. The claim 3 then follows from the connectedness of $S$. Given any point $x$ in $S$, there is a path from $x$ to some point in $T$. This path traverses a finite number of neighbouring tiles, including any tile containing $x$. Hence any tile containing $x$ can be taken to be one of the $T_{i_{j}}$ above, and is mapped by $\psi_{i}$ to some other tile of $S$.
@@ 4. Finite edge type oriented tiles
where, if we write $e\equiv e^{\prime}$ to denote that there is a rigid motion mapping the edge $e$ to the edge $e^{\prime}$, preserving the surface normal along the edge, we have:
@@ 4.1. Smoothability of monohedral polyhedra
Let $\Sigma$ be a compact finite edge type tiled surface with prototile of boundary configuration $\alpha\alpha\beta$ (Figure 6). Suppose that $\Sigma$ has vertices of degrees of both degree $4$ and degree $5$. Denote the interior angles at the corners $A$,$B$ and $C$ of $T$ by $\mathbf{\theta_{a}}$, $\mathbf{\theta_{b}}$ and $\mathbf{\theta_{c}}$. Then the interior angles must satisfy one of the following pairs of equations:
Case 1: $\left\{\begin{array}[]{rll}5\mathbf{\theta_{a}}&=2\pi&(\alpha\alpha\alpha\alpha\alpha)\\ 2\mathbf{\theta_{b}}+2\mathbf{\theta_{c}}&=2\pi&(\alpha\beta\alpha\beta)\end{array}\right.$
Case 2: $\left\{\begin{array}[]{rll}5\mathbf{\theta_{a}}&=2\pi&(\alpha\alpha\alpha\alpha\alpha)\\ \mathbf{\theta_{b}}+\mathbf{\theta_{c}}+2\mathbf{\theta_{a}}&=2\pi&(\alpha\beta\alpha\alpha)\end{array}\right.$
Case 3: $\left\{\begin{array}[]{rll}2\mathbf{\theta_{b}}+2\mathbf{\theta_{c}}+\mathbf{\theta_{a}}&=2\pi&(\alpha\beta\alpha\beta\alpha)\\ 2\mathbf{\theta_{a}}+\mathbf{\theta_{b}}+\mathbf{\theta_{c}}&=2\pi&(\alpha\beta\alpha\alpha)\end{array}\right.$
Case 4: $\left\{\begin{array}[]{rll}2\mathbf{\theta_{b}}+2\mathbf{\theta_{c}}+\mathbf{\theta_{a}}&=2\pi&(\alpha\beta\alpha\beta\alpha)\\ 4\mathbf{\theta_{a}}&=2\pi&(\alpha\beta\alpha\alpha)\end{array}\right.$
Case 5: $\left\{\begin{array}[]{rll}\mathbf{\theta_{b}}+\mathbf{\theta_{c}}+3\mathbf{\theta_{a}}&=2\pi&(\alpha\beta\alpha\alpha\alpha)\\ 2\mathbf{\theta_{b}}+2\mathbf{\theta_{c}}&=2\pi&(\alpha\beta\alpha\beta)\hbox{or }(\alpha\alpha\alpha\alpha)\end{array}\right.$
Cases $1$ to $4$ have only one type of vertex configuration of degree $5$ and one of degree $4$. Case $5$ has only the vertex configuration $(\beta\alpha\beta\alpha\alpha)$ of degree $5$ but may have either or both vertex configurations $(\alpha\beta\alpha\beta)$ and $(\alpha\alpha\alpha\alpha)$ of degree $4$
@@
(7) $5\mathbf{\theta_{a}}=2\pi,\quad\hbox{or}\quad 2\mathbf{\theta_{b}}+2\mathbf{\theta_{c}}+\mathbf{\theta_{a}}=2\pi,\quad\hbox{or}\quad\mathbf{\theta_{b}}+\mathbf{\theta_{c}}+3\mathbf{\theta_{a}}=2\pi$
@@ 5.1. Prototiles with $1$ or $2$ edges
The only configurations possible for a tile with two edges are $a\bar{a}$ and $\alpha\alpha$. For $a\bar{a}$, if we denote the two vertices by $p$ and $q$, and noting that $\bar{a}$, which runs from $p$ to $q$ is reverse congruent to $a$, which runs from $q$ to $p$, it follows that the congruence is given by a rotation about the axis through $p$ and $q$. Since the interior of the tile is by definition non-empty, this rotation is by some nonzero angle $\theta$. There is only one way for the tiling to continue, and that is to repeatedly rotate the prototile by the same angle $\theta$ about this axis. If we require the surface to be embedded, the angle must be of the form $\theta=2\pi/m$, where $m$ is an integer, and the surface is a tiling of a topological sphere by $m$ copies of this tile. We also have a restriction that the tangent to the edge curve $a$ at the end points must be perpendicular to the axis, and that the tangent to the tile at the vertices is the plane perpendicular to the axis. The set of such surfaces (if embedded) is the set of compact surfaces of genus zero that are invariant under a rotation of finite order (Figure 7). For $\alpha\bar{\alpha}$, one also has the possibility to turn the tile around, reversing $p$ and $q$, which gives various shapes depending on the interior of the tile, but the result is still a topological sphere.

### v2

@@ Abstract.
We study $C^{1}$-regular surfaces in $\mathbb{R}^{3}$ that admit tilings by a finite number of rigid motion congruence classes of tiles. We construct examples with various topologies and present a framework for a systematic study, mainly concentrating on monotilings. A finite edge prototile is a tile that has only a finite number of possible interfaces with adjacent copies of itself. We describe all monotilings by such tiles with three or less edges. We consider the question of whether a monohedral polyhedron can be “smoothed” to become a finite edge type tileable surface with the same graph structure, and we give an example where this is not possible. Finally we list some open problems.
@@
Research partly supported by Horizon-EIC-2023 pathfinder challenge 01 project number 101161085 STACK
@@ 1. Introduction
Tilings of the plane, sphere and hyperbolic space are well known. But if we consider physical manifestations of tiling, that is, a figure or object composed of many copies of the same shape, or a limited number of distinct shapes, there is no special reason why a given tile should be of constant curvature, or that the total space need be homogeneous (i.e., possessing global isometries mapping any point to any other point). In manufacturing and architectural scenarios, there is an economic case for constructing a surface from many copies of a limited number of distinct shapes or tiles. Recently, there has been some serious work related to this problem, in the special case of polyhedral surfaces and variants – see the literature review in Section 2 below.
@@
where $m\in{\mathbb{N}}\cup\{\infty\}$ is an extended natural number and,
@@
Even if we allow for a large (finite) number $n$ of prototiles, it is easy to see that there are restrictions on the global shape of a tileable surface. For instance (see, e.g. [docarmo1, lee2003smooth] for terminology):
@@
To see that this holds, note that, since $S$ is tiled by a finite number of distinct compact prototiles, the curvature must achieve both its maximum and minimum values, and hence be bounded away from zero. By Efimov’s theorem [efimov1968hyperbolic], there is no complete $C^{2}$-immersed surface in $\mathbb{R}^{3}$ with curvature bounded above by a negative constant. If the curvature is positive the surface must be compact by the Bonnet-Myers theorem, and hence a topological sphere by the Gauss-Bonnet theorem.
Similarly, given a $C^{2}$ monotiled surface, although it is obviously possible to deform the prototile and apply the corresponding deformation to all the other copies to remain a $C^{2}$ monotiled surface with the same topology and face combinatorics, there are nevertheless restrictions on the deformed shape. By the Gauss-Bonnet theorem we have:
@@
    Tileable surfaces with boundary.
@@ 1.1. Relation between tileable surfaces and tilings of the plane or sphere
Figure 2. From left to right: monohedral polyhedron (biarc-hull [eppstein2021polyhedral]); projection to the sphere; more general $C^{1}$ tiling; geometrically distinct tiled surface with the same prototile.
@@ 1.2. Relation between tileable surfaces and monohedral polyhedra
One idea for producing examples of tileable surfaces is to take a monohedral polyhedron and, if possible, deform the model face until the faces fit together tangent continuously along the edges (Figure 2). This is obviously possible in some cases. However, we will show that there are also examples that cannot be smoothed in this way (See Theorem 12 below and Figure 3). Thus, a basic problem is to understand which monohedral polyhedra can be smoothed to become regular tiled surfaces.
Conversely, given a tiled surface, if we define an “edge” on the boundary of a tile to be an uninterrupted interface between two tiles, and call its endpoints “vertices”, then the vertices may correspond to those of a polyhedral surface. This can be done for the first surface in Figure 1, to form a tiled plane. For the torus shown in the same figure, the resulting polyhedron collapses because the vertices are all in the same plane; more generally, if the number of vertices on a face is more than three, they will not usually be co-planar, and moreover for some tiled surfaces (Figure 4) there are no canonical points to choose as vertices on the tile.
@@ 2.1. Doubly and triply periodic surfaces
These are true tileable surfaces, according to our definition. Any 2-dimensional or 3-dimensional lattice structure gives us many examples of monotiled surfaces, simply by placing a surface in the fundamental domain in such a way that translated copies of it are joined tangent continuously. A deformation of a periodic plane-tiling (Figure 1, left) is an example of a doubly periodic surface. More generally, so-called triply periodic surfaces in $\mathbb{R}^{3}$ can easily be constructed by placing a suitable surface inside the fundamental domain of a 3-dimensional lattice. It is possible to arrange that the fundamental piece is also tiled by congruent copies of a topological disc, thus the entire surface satisfies our definition of a monotiled surface. Triply periodic minimal surfaces, first discovered by Schwarz and his student in the 19th century, are well known (see, e.g. [meeks1990theory], [karcher1996construction], [grosse2012triply]) and have recently become very popular in practical applications such as in additive manufacturing and in tissue engineering: see, e.g., [rajagopalan2006schwarz], [feng2022triply] and many other recent works.
@@ 2.2. Tilings of constant curvature surfaces
This topic is well known, and extensively studied. See [grunbaum1987tilings], [schulte1993tilings], [toth2017handbook] and [adams2023tiling], for an introduction to tilings, and [zong2020can] for a survey on monotiling of the plane. Common themes are classifications of patterns, group actions, tilings with as small a number of distinct tiles as possible, non-repeating tilings, tilings with different edge colorings etc. However, we have been unable to find any systematic attempt to extend this study to smooth surfaces of non-constant curvature.
@@ 2.3. Polyhedral surfaces
It follows from a theorem of Steinitz that every 3-connected planar graph can be realized as a convex polyhedron. However, there are restrictions on the geometry: for instance, not every such graph can be realized as a monohedral polyhedron. Grünbaum [grunbaum2001convex] shows that there exist triangulated spheres that are not realized by any acoptic (meaning homeomorphic to a sphere) monohedral polyhedron.
A monohedral polyhedron is called an isohedron if all faces lie within the polyhedron’s symmetry orbit, i.e., the rigid motions of the entire polyhedron act transitively on the faces. Examples include the Platonic solids, the Catalan solids, the bipyramids and the trapezohedra. Convex isohedra are classified, with a list of 30 types. Even in the convex case, general monohedral polyhedra do not appear to be classified. Examples that are not isohedra include the triaugmented triangular prism (Figure 3), which has equilateral triangular faces, and the rhombic icosahedron which has 20 rhombic faces. Eppstein [eppstein2021polyhedral] gives an infinite family of monohedral polyhedra monotiled by isosceles triangles.
@@ 2.4. Polyhedral and semidiscrete approximations
Eigensatz et al [eigensatz2010paneling] describe the “paneling problem” where a free form surface is to be approximated by as few as possible distinct panels to minimize the number of molds to be used in the construction. Liu et al [liu2023reducing] study a similar problem, using polygonal faces. Singh and Schaeffer [singh2010triangle] give an algorithm where the input is an arbitrary triangulated surface and the output is an approximation using a small number of polygonal faces. Fu et al [fu2010k] give an algorithm for optimization of a quadmesh approximation of a surface using a small number of distinct (non-planar) quadrilaterals.
Chen et al [chen2023masonry] consider a similar problem for a shell structure: approximate the surface by a piecewise linear shell structure, where the goal is to have many identical parts. The difference from polyhedral surfaces is the thickness.
Bo et al [bo2011circular] (see also [bartovn2013circular]), replace polyhedral surfaces with circular arc structures, that is surfaces defined by 2D mesh combinatorics, where the edges are circular arcs that meet with a common tangent plane. Their study shares features of what we discuss below, in that the edges of the faces are tangent continuous. However they do not consider surfaces with congruent faces - instead they aim to construct surfaces with congruent vertices.
@@ 2.5. Mosaic methods
Another approximation concept for smooth surfaces, mosaic methods [PW2008], [PW2009] [chen2017fabricable],[hu2015surface], means approximating a 3D surface by covering as much as possible of it with flat tiles drawn from a given collection. Small gaps are allowed between the tiles. The measure of success for a tile set together with an algorithm is to minimize the gap area while closely approximating the surface.
@@ 2.6. Tiling inspired decoration of arbitrary surfaces
Any surface has the conformal structure of one of the standard 3 constant curvature space forms, which allows for texture maps conformally equivalent to the many known tilings of space forms (see, e.g. Gu et al [gu2010fundamentals]). These are not tilings in the sense discussed here, however, since conformal maps preserve angles but not size.
Jiang et al [jiang2015polyhedral] explore the idea of extending the common tiling patterns on the plane to polyhedral surfaces by another idea. The surface tiles are defined by lifting planar tiles onto paraboloids (approximating the surface around a point by the curvature paraboloid). The method gives a natural way to produce tiling-like patterns on general surfaces, but the tiles are not congruent to each other. The shape of the lift to the surface of a particular model tile will depend on the curvature of the surface at that point.
@@ 3. Rigid tilings
    There is a unique set $\{T_{1},T_{2},\dots,T_{N}\}$ that constitutes the set of neighbours of $T$ in any tiled surface containing $T$ as an interior tile.
@@
The tiles in Figure 1 are rigid, whilst Figures 2 and 4 show examples of tiles that are not rigid. The tile in Figure 2 can be surrounded by itself in only a finite number of ways. The tile in Figure 4 has a circle as its boundary, and surface normal on the boundary parallel to the radius of the circle. There are uncountably many ways to fit two copies together.
@@
    and that each set $\psi_{i}(T_{i_{j}})$ is in fact a tile of $S$. The claim 3 then follows from the connectedness of $S$ (see Definition 1). Given any point $x$ in $S$, there is a path from $x$ to some point in $T$. This path traverses a finite number of neighbouring tiles, including any tile containing $x$. Hence any tile containing $x$ can be taken to be one of the $T_{i_{j}}$ above, and is mapped by $\psi_{i}$ to some other tile of $S$.
@@ 4. Finite edge type oriented tiles
where, if we write $e_{1}\equiv e_{2}$ to denote that there is a rigid motion mapping the edge $e_{1}$ to the edge $e_{2}$, preserving the surface normal along the edge, we have:
@@ 4.1. Smoothability of monohedral polyhedra
Let $\Sigma$ be a compact finite edge type tiled surface with prototile of boundary configuration $\alpha\alpha\beta$ (Figure 6). Suppose that $\Sigma$ has vertices of both degree $4$ and degree $5$. Denote the interior angles at the corners AA,$B$ and $C$ of $T$ by $\mathbf{\theta_{a}}$, $\mathbf{\theta_{b}}$ and $\mathbf{\theta_{c}}$. Then the interior angles must satisfy one of the following pairs of equations:
Case 1: $\left\{\begin{array}[]{rll}5\mathbf{\theta_{a}}&=2\pi,\quad\quad&(\alpha\alpha\alpha\alpha\alpha)\\ 2\mathbf{\theta_{b}}+2\mathbf{\theta_{c}}&=2\pi,\quad\quad&(\alpha\beta\alpha\beta)\end{array}\right$.
Case 2: $\left\{\begin{array}[]{rll}5\mathbf{\theta_{a}}&=2\pi,\quad\quad&(\alpha\alpha\alpha\alpha\alpha)\\ \mathbf{\theta_{b}}+\mathbf{\theta_{c}}+2\mathbf{\theta_{a}}&=2\pi,&(\alpha\beta\alpha\alpha)\end{array}\right$.
Case 3: $\left\{\begin{array}[]{rll}2\mathbf{\theta_{b}}+2\mathbf{\theta_{c}}+\mathbf{\theta_{a}}&=2\pi,\quad\quad&(\alpha\beta\alpha\beta\alpha)\\ 2\mathbf{\theta_{a}}+\mathbf{\theta_{b}}+\mathbf{\theta_{c}}&=2\pi,\quad\quad&(\alpha\beta\alpha\alpha)\end{array}\right$.
Case 4: $\left\{\begin{array}[]{rll}2\mathbf{\theta_{b}}+2\mathbf{\theta_{c}}+\mathbf{\theta_{a}}&=2\pi,\quad\quad&(\alpha\beta\alpha\beta\alpha)\\ 4\mathbf{\theta_{a}}&=2\pi,\quad\quad&(\alpha\beta\alpha\alpha)\end{array}\right$.
Case 5: $\left\{\begin{array}[]{rll}\mathbf{\theta_{b}}+\mathbf{\theta_{c}}+3\mathbf{\theta_{a}}&=2\pi,\quad\quad&(\alpha\beta\alpha\alpha\alpha)\\ 2\mathbf{\theta_{b}}+2\mathbf{\theta_{c}}&=2\pi,\quad\quad&(\alpha\beta\alpha\beta)\hbox{ or }(\alpha\alpha\alpha\alpha)\end{array}\right$.
Cases $1$ to $4$ have only one type of vertex configuration of degree $5$ and one of degree $4$. Case $5$ has only the vertex configuration $(\beta\alpha\beta\alpha\alpha)$ of degree $5$ but may have either or both vertex configurations $(\alpha\beta\alpha\beta)$ and $(\alpha\alpha\alpha\alpha)$ of degree 44
@@
(7) $5\mathbf{\theta_{a}}=2\pi,\quad\hbox{or}\quad 2\mathbf{\theta_{b}}+2\mathbf{\theta_{c}}+\mathbf{\theta_{a}}=2\pi,\quad\hbox{or}\quad\mathbf{\theta_{b}}+\mathbf{\theta_{c}}+3\mathbf{\theta_{a}}=2\pi,$
@@ 5.1. Prototiles with $1$ or $2$ edges
The only configurations possible for a tile with two edges are $a\bar{a}$ and $\alpha\alpha$. For $a\bar{a}$, if we denote the two vertices by $p$ and $q$, and noting that $\bar{a}$, which runs from $p$ to $q$ is reverse congruent to $a$, which runs from $q$ to $p$, it follows that the congruence is given by a rotation about the axis through $p$ and $q$. Since the interior of the tile is by definition non-empty, this rotation is by some nonzero angle $\theta$. There is only one way for the tiling to continue, and that is to repeatedly rotate the prototile by the same angle $\theta$ about this axis. If we require the surface to be embedded, the angle must be of the form $\theta=2\pi/m$, where $m$ is an integer, and the surface is a tiling of a topological sphere by $m$ copies of this tile. We also have a restriction that the tangent to the edge curve $a$ at the end points must be perpendicular to the axis, and that the tangent to the tile at the vertices is the plane perpendicular to the axis. The set of such surfaces (if embedded) is the set of compact surfaces of genus zero that are invariant under a rotation of finite order (Figure 7). For $\alpha\alpha$, one also has the possibility to turn the tile around, reversing $p$ and $q$, which gives various shapes depending on the interior of the tile, but the result is still a topological sphere.